OpenStudy (anonymous):

-solve using only elimination for L and W. Please Help. L×W=1200 π×(W/2π)2×L=600

OpenStudy (anonymous):

If you take the first equation and solve for W, what is W equal to (in terms of L)?

OpenStudy (anonymous):

Then you can substitute that value for W into the second equation.

OpenStudy (anonymous):

i have to use elimination not substitution for the first part :/

Directrix (directrix):

@pmanrique --> Is ×(W/2π)2 to be read as ×(W/2π)^2 or just times two?

OpenStudy (anonymous):

its an exponent so the second one sorry

OpenStudy (anonymous):

but only using elimination method not substitution at least for the beginning

OpenStudy (anonymous):

Sorry -- so with elimination, you "add" the two equations together -- add the left side of the first to the left side of the second, then the right side of the first to the right side of the second. What does that get you?

OpenStudy (anonymous):

i don't even know :(

OpenStudy (anonymous):

Sorry -- instead of adding the equations, try dividing the second equation by the first. The right side of the equation would look like: \[600/1200\] -- what would the left side look like?

OpenStudy (anonymous):

well i did this: L*(Pi)(w/2Pi)^2=600 -(L*W)=1200 so then the L's would cancel but now i can't figure out how to do the w's

OpenStudy (anonymous):

With elimination, when you divide the equations, you write a third new equation. \[(L*\Pi*(w/2\Pi)^2)/(L*w) = 600/1200\] With that, the L's will fall out and leave you with W.

Directrix (directrix):

(1) π×(W/2π)^2×L=600 (2) LW = 1200 --------------------- The first equation's left side simplifies to π³ /4 L W². So, π³ /4 L W² = 600 Multiply the second equation by - π³ /4 W That gives - π³ /4 W (LW) = (- π³ /4 W) 1200 So, the system to be solved by elimination becomes: π³ /4 L W² = 600 - π³ /4 W (LW) = (- π³ /4 W) 1200 ----------------------------------------- π³ /4 L W² - π³ /4 W (LW) = 600 + (- π³ /4 W) 1200 0 = 600 + (- π³ /4 W) 1200 -600 = (- π³ /4 W) 1200 1/2 = π³ /4 W 2/ π³ = W Now, solve for L --> Check my work so far.